A rocket is fired ,with an acceleration of 20m/sec2 for the first 1min.After its fuel is finished and it goes on rising up just like a  free body .Up to which height will it rise?Also find its time it will take before falling on earth?(g=10m/sec2)

Grade:11

1 Answers

Arun
6 years ago
Velocity for 20m/s² accelration
V1= u + at
= 0 + 20 x 1 60 
=1200 m/s

Height = h = ut + 1/2 + 20 + 60²
= 36000m

Velocity for 10 m/s²
v2 = v1 + gt'
0 = 1200 - 10t
t' = 120 s

Height for 10 m/s²

h'= v1t' + 1/2gt'²
= 1200 x 120 -1/2 x 10 x 120²
= 72000m

total height = h+h'
= 36000 + 72000
= 108000m

Time taken 
h=1/2 gt²
t = √2hg
= √2 x 108000 x 10
= 600√6 sec or =60 (3+√6)